OpenStudy (anonymous):

graph each system of constraints. Name all vertices. Then find the values of x and y that maximize or minimize the objective function. {6<= x + y <= 13 {x >=3 {y >=1 Maximum for P=4x + 3y

jimthompson5910 (jim_thompson5910):

Do you know how to graph these inequalities?

OpenStudy (anonymous):

No I have no clue :-(

jimthompson5910 (jim_thompson5910):

well the first one is a bit strange, but it means that the line y = -x is between 6 and 13 on the y axis (basically have the line y = -x+6 and y = -x+13 and shade everything between them)

jimthompson5910 (jim_thompson5910):

x >= 3 means you have a vertical line at x = 3 and you shade to the right

jimthompson5910 (jim_thompson5910):

y >= 1 means you have a horizontal line at y = 1 and you shade above this line

OpenStudy (anonymous):

ok That makes sense so far.

jimthompson5910 (jim_thompson5910):

all 4 regions will overlap to form a new region

jimthompson5910 (jim_thompson5910):

the corners of this region are potential max/min points so you must find each corner point and you must test each one in P=4x + 3y to see which point is the max

OpenStudy (anonymous):

how do I find each corner point?

jimthompson5910 (jim_thompson5910):

each corner is created by 2 lines intersecting

jimthompson5910 (jim_thompson5910):

so you must solve a system of equations (2 equations 2 unknowns) to find each corner point

jimthompson5910 (jim_thompson5910):

for instance, solving the system -x - y = -6 y = 1 gives you the corner point (5,1)

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

you should get 4 corner (or vertex) points

jimthompson5910 (jim_thompson5910):

tell me which vertex points you get

OpenStudy (anonymous):

ok still trying to figure it out

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

would one of the corners be (3,1)

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

that's one intersection, but it's not in the shaded region

OpenStudy (anonymous):

(3, -9)

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

here's what the shaded region looks like

OpenStudy (anonymous):

I dont know why I have such a problem grasping this. I have never been good at graphs and such

jimthompson5910 (jim_thompson5910):

you'll get better with more practice

jimthompson5910 (jim_thompson5910):

So from this graph, we see that the vertex points are (3, 3), (5, 1), (3, 10), (12, 1)

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

From here, plug each vertex point into the objective function P = 4x+3y to find the min and max Plug in (3,3) P = 4x+3y P = 4(3)+3(3) P = 21 --------------------------------------------- Plug in (5, 1) P = 4x+3y P = 4(5)+3(1) P = 23 --------------------------------------------- Plug in (3,10) P = 4x+3y P = 4(3)+3(10) P = 42 --------------------------------------------- Plug in (12, 1) P = 4x+3y P = 4(12)+3(1) P = 51 --------------------------------------------- Summary Min is P = 21 which occurs at (3,3) Max is P = 51 which occurs at (12,1)

OpenStudy (anonymous):

Ohh okay. That part makes sense! Thank you so much! Now I just hope I can get that graphing part down for good, :-)

jimthompson5910 (jim_thompson5910):

you will, just spend more time doing it and you'll get better

jimthompson5910 (jim_thompson5910):

I would recommend practicing graphing linear equations. The inequality part is just the shading (and how the line is drawn -- either solid or dashed).

OpenStudy (anonymous):

Okay. I will definitely do that. Thanks for all your help.

jimthompson5910 (jim_thompson5910):

yw